OpenStudy (anonymous):

Algebra questions:

OpenStudy (anonymous):

@misty1212 \[what is the value of -3x ^{2}y ^{3} when x = 2 and y = -1?\]

OpenStudy (misty1212):

\[3\times 2^2\times (-1)^3\] is a start

OpenStudy (misty1212):

well actually if i read it right \[-3\times 2^2\times (-1)^3\]

OpenStudy (misty1212):

you have 4 minus signs so it is positive, then compute \[3\times 2\times 2\]

OpenStudy (anonymous):

= 12?

OpenStudy (bee_see):

12 is correct.

OpenStudy (anonymous):

Thanks. Will you help me with another question? @Bee_see

OpenStudy (bee_see):

ok

OpenStudy (anonymous):

Draw a rectangle that meets both of the following conditions and label the rectangle's length and width: the perimeter of the rectangle is 36 units. the length of the rectangle is 5 times its width.

OpenStudy (bee_see):

solve for x. add like terms

OpenStudy (bee_see):



OpenStudy (bee_see):

in this case, you just add the numbers in front of x. For example: 2x+3x=5x

OpenStudy (bee_see):

add 5x+5x+1x+1x

OpenStudy (bee_see):

5+5+1+1?

OpenStudy (anonymous):

they both equal 12x

OpenStudy (bee_see):

Yes. now we have 36=12x

OpenStudy (bee_see):

solve for x. what's the the opposite of multiplication?

OpenStudy (bee_see):